# Improving probabilistic forecasts of extreme wind speeds by training statistical post-processing models with weighted scoring rules

### v2

@@ 4 Weighted scoring rules for assessing forecasts of extremes
The goal of probabilistic forecasting has been put by Gneiting et al. (2005) as to "maximize sharpness subject to calibration". Sharpness is a measure of the forecast distribution only and refers to the concentration of the forecast density function, whilst calibration is a joint property between observations and forecast and corresponds to the consistency between the forecast distributions and the observations. Proper scoring rules (Gneiting and Raftery, 2007) have been introduced for the evaluation (and training) of probabilistic forecasts. They assess both calibration and sharpness. Briefly, a scoring rule $S$ is proper when the following holds for any probabilistic forecast $F$ and "true" distribution $G$:
@@
Here $S(F,y)$ is the score of the predictive CDF at the observation $y$ and $\mathbb{E}_{y\sim G}S(F,y)$ denotes its expectation over the distribution of $G$. A scoring rule is thus called proper when in expectation it is minimal at the true distribution. It is called strictly proper if equality in eq. 7 holds only when $F=G$.
@@
where $g$ is the density corresponding to $G$. Thus when evaluating probabilistic forecasting models using scoring rules restricted to an outcome (such as events over a threshold) or scoring rules weighted by some weighting function (such as the HB-loss) this encourages hedging of the forecast distribution, meaning a forecaster is encouraged to deviate from the true distribution in their prediction by an amount determined by the weighting function. This also affects model training as can be seen in section 3.
@@
for a nonnegative weighting function $w$. This reduces to the unweighted CRPS (eq. 4) when $w(z)\equiv 1$. The canonical weighting is indicator weighting: $w_{\tau}(z)=\mathbb{1}\{z\geq\tau\}$, when there is interest in values above (or below) a certain threshold $\tau$, which then leads to the so-called threshold-weighted CRPS (twCRPS):
@@
The weighted CRPS and especially the threshold-weighted versions have become frequently used tools for forecast evaluation of extreme events (e.g. Lerch and Thorarinsdottir, 2013; Baran and Lerch, 2015; Baran and Nemoda, 2016; Allen et al., 2021b). They have also been generalized to multivariate versions (Allen et al., 2023b, a). In the following we will focus on the twCRPS as given in equation 10.
@@
This is alluded to already in Allen et al. (2023a). The proof is given in the appendix. Thus intuitively when using the twCRPS a forecaster effectively uses the full CRPS to evaluate the distribution $\tilde{F}_{\tau}$ that has all probability mass below the threshold collapsed onto the threshold itself, whilst remaining the same above the threshold. That way the exact form of the distribution below $\tau$ only influences the score via the probability mass for exceeding or falling below the threshold.
@@
where $v_{\tau}$ is the so-called "chaining function":
$v_{\tau}(x):=\begin{cases}x_{0}&\text{ if }x<\tau\\ x&\text{ if }x\geq\tau\end{cases}.$ (14)
Here $x_{0}$ is an arbitrary point that all values below the threshold $\tau$ are mapped to. In the following we will use the canonical choice $x_{0}=\tau$ for which the chaining function is continuous. We will discuss the impact of other choices of $x_{0}$ on sample based optimization of the twCRPS in appendix D. The formulation of the twCRPS as an expectation can be used to approximate it via samples from a predictive distribution, i.e. using the estimator
@@
Although the twCRPS has been frequently used for evaluation of probabilistic forecasting models for extremes, it has – to the authors’ best knowledge – to this date not been used for training post-processing models i.e. as a loss function for parameter estimation. However, given that weighted and threshold-weighted scores are proper – something that section 3 showed is important also for model training – and the relevance of reliable forecasts for extremes, we will study their usage for the training of EMOS post-processing models for wind speed, for both truncated normal and truncated logistic EMOS models.
@@ 5 Results: training with weighted scoring rules
We study the usage of the weighted CRPS for the training of post-processing models, to improve probabilistic predictions of extreme wind speeds. The weighted CRPS is a proper scoring rule, thus using it for evaluation and training will favour calibrated-sharp forecasts. We focus on the canonical threshold weighting due to the availability of closed-form solutions, which helps with the optimization. We consider training EMOS models using the following proper scores:
@@
Here the CRPS and log-likelihood models serve as baselines, whilst the twCRPS puts special emphasis on values above a threshold $\tau$. For optimizing the score functions we use the closed-form expressions for the CRPS and twCRPS for the truncated normal and truncated logistic distributions (see appendix B).
@@ 5.1 Effect of weighted training
We consider thresholds $\tau$ defined as the location-wise 80th and 90th percentile of the training dataset and train using the scores in section 5. In practical applications most likely absolute thresholds defined by user needs will be of relevance, however to show the general applicability of the method we use percentile based thresholds here as in Sharpe et al. (2018); Sharpe and Stewart (2019); Sharpe (2022). Figure 12 in the appendix shows the 80th and 90th percentile threshold at each station. We report skill scores relative to the model trained using CRPS ("CRPS-trained model")
(a) Evaluated using twCRPSS
(b) Evaluated using CRPSS
@@
Figure 3: Threshold-weighted CRPS skill score (left, a) and CRPS skill score (right, b), both relative to the CRPS-trained model, for wind speed EMOS using truncated normal and truncated logistic distributions. Models are trained using log likelihood, and twCRPS. Thresholds for training and evaluation of the twCRPS are defined as locationwise 80th and 90th percentile of the training dataset.
@@
Figure 3(a) shows for all models the threshold-weighted CRPS skill score over the CRPS-trained model evaluated on the test set, for the 80th and 90th locationwise percentile thresholds (used for both training and evaluation) for both truncated normal and truncated logistic predictive distributions. Training using a twCRPS leads to improved predictions of extremes, as measured by the twCRPS on the test set. These improvements average around $\sim 2.5$% for the 80th and around 5% for the 90th percentile threshold and up to $\sim 8$% and $\sim 15$% respectively. The twCRPS improvements also show in terms of Brier skill score where we see that by training using a twCRPS one also gets better predictions of the occurrence of threshold exceedances of around 1.7% for the 80th and around 0.9% for the 90th percentile threshold in the truncated normal and around 1.6% and 0.7% in the truncated logistic case (see table 1 in the appendix for the Brier skill score for both distributions for a variety of evaluation thresholds). Finally the improvements seem to be distributed heterogeneously over space, possibly with slightly stronger improvements for stations closer to the coast (Figure 13 in the appendix). The log-likelihood training performs worse in terms of CRPS and twCRPS than the CRPS-trained model which is in line with previous studies (Gneiting et al., 2005; Gebetsberger et al., 2018).
@@
The twCRPS improvement seems however to be linked to a loss in CRPS as can be seen in Figure 3(b) which shows the CRPS skill over the CRPS-trained model, evaluated on the test set. The twCRPS-trained model loses CRPS skill, compared to the CRPS-trained model for both the 80th and even stronger for the 90th percentile threshold. The latter threshold improves more in terms of twCRPS, but also loses more in terms of CRPS, irrespective of the predictive distribution.
There seems to be a trade-off between CRPS and twCRPS performance, which we investigate more closely. We focus on the truncated normal distribution, the 80th percentile and twCRPS training. Figure 4(a) shows the twCRPS skill improvement against the CRPS skill loss for all locations. We see a trade-off between the two objectives for this model, which have different ranges of skill change, meaning that a twCRPS improvement might lead to a much stronger CRPS loss. Figure 4(b) shows that the CRPS loss and twCRPS improvement relate to the distance between the EMOS parameters of the CRPS-trained and twCRPS-trained model. The figure shows the Euclidean norm of the difference between the EMOS parameters of the twCRPS-trained and CRPS-trained model, plotted against CRPS skill loss (blue) and twCRPS skill improvement (orange), both relative to the CRPS-trained model. The bigger the difference in EMOS parameters, the bigger the twCRPS improvement can be, however this comes with a larger CRPS skill loss as well.
(a) CRPS skill loss against twCRPS skill gain
(b) Parameter difference against twCRPS/CRPS change
@@
Figure 4: CRPS-twCRPS trade-off for the truncated normal, twCRPS-trained model for the 80th percentile. Left (a): CRPS skill loss against twCRPS skill gain for all locations, both relative to the CRPS-trained model. Right (b) Difference in EMOS parameters between the twCRPS-trained model and the CRPS-trained model plotted against CRPS (blue) or twCRPS (orange) skill over the CRPS-trained model
@@
In summary, by optimizing the twCRPS one can improve extreme event performance of forecasts, as measured by twCRPS and Brier Score. This however comes at some cost in terms of CRPS – measuring the forecast performance for the whole distribution and the distribution body. The trade-off between CRPS and twCRPS performance can be seen as a classic distribution "body-tail trade-off" where an improved modelling of a distribution’s tails might come at the cost of capturing of the distribution’s body and vice-versa. This seems to hold for training thresholds defined as 80th and 90th percentile of the distribution, however in the next section we will study the effect of varying the training threshold.
@@ 5.2 Effect of the training threshold
We study the effect of varying the threshold for training using the twCRPS and analyze this for the truncated normal case. Results for the truncated logistic distribution are similar. We train post-processing models using the twCRPS with thresholds as 60th, 80th, 90th and 95th percentile and evaluate against a range of different thresholds between the 0th (corresponding to the unweighted CRPS) to 99th percentile of the locationwise training dataset. Figure 5 shows the twCRPS skill (relative to the CRPS-trained model) as a function of the threshold for all four trained models. The model trained for a certain threshold seems to perform best at this threshold (and slightly lower) and has generally positive skill score for evaluation thresholds higher than the one it was trained for. However training with a twCRPS for a high quantile threshold leads to skill loss at lower quantiles and it seems the higher the training threshold / quantile the worse the twCRPS skill loss for lower evaluation thresholds / quantiles also becomes, manifesting in increasing slope of the skill curves. This means that model performance can be improved for many thresholds by training with a twCRPS, but should the distribution body performance be important it might be sensible to use lower thresholds. In these cases however the skill improvements for high quantiles also do not seem to be as substantial. Similar curves are also obtained when evaluating the Brier score (see figure 14 in the appendix).
@@
Figure 5: twCRPS skill score (over the CRPS-trained model) as a function of the evaluation threshold for twCRPS-trained models trained using different thresholds. The bands represent 95% intervals over stations. The twCRPS for an evaluation quantile of $0.0$ corresponds to the unweighted CRPS.
5.3 Effect of sample-based estimation
In the sections above we have investigated the impact of training post-processing models using a twCRPS, optimizing closed-form solutions for the twCRPS for both truncated normal and truncated logistic distributions. However closed-form solutions like these might not be available in all cases, for example when a model does not follow a known distributional structure. In these cases the predictive distribution function might still be available via samples and one can estimate the twCRPS by approximating its representation as an expectation (equation 13). In the following we study whether sample based optimization of the twCRPS can also be effective for training models. We study again EMOS using the truncated normal distribution, trained using the twCRPS with a location-wise 80th and 90th percentile threshold. However we only compute the twCRPS by sampling from the model.
We find that sample-based optimization proves difficult. First the optimization algorithm has a strong influence. The BFGS algorithm frequently fails to find optimal parameters, even when using 1000 samples for each predictive distribution, which is most likely due to noisy empirical gradients and a noisy and relatively flat cost function. The Nelder-Mead optimizer, which is not based on gradients, can converge in some cases and for some starting values, after adjusting the tolerance to account for fluctuations in the twCRPS values. Similarly Adam (Kingma and Ba, 2017), a stochastic gradient descent algorithm, together with the reparametrization trick (Kingma and Welling 2022, see appendix C) to allow gradients with respect to the parameters, can converge with the right starting values. However, it is much slower than Nelder-Mead or BFGS-based optimization of the closed-form solutions. It also leads to EMOS models which are much less efficient i.e. have far smaller twCRPS improvement than training using the closed-form expression, if they have any improvement at all.
The sample-based twCRPS seems to represent a rather noisy and possibly very flat loss objective, whose optimization seems dominated by sampling uncertainty. In certain cases (e.g. with starting values that are close to the solution and using exact gradients) the optimization can converge, however it is quite unstable. The results speak for using closed-form expressions for the optimization in cases in which these are available (see appendix B for a list of closed-form expressions) as the optimization objective seems much simpler to solve. However closed-form expressions might not be available in all cases, e.g. for very complex models, other weighting functions different from threshold weighting or when wanting to use multivariate weighted scores (Allen et al., 2023b, a). We defer the question of how to stabilize sample-based twCRPS optimization in these cases for future work. Also, the interested reader can find more analysis on sample based optimization in appendix D.

### v4

@@ 4 Weighted scoring rules for assessing forecasts of extremes
Proper scoring rules (Gneiting and Raftery, 2007) have been introduced for the evaluation (and training) of probabilistic forecasts. They assess both calibration and sharpness. Briefly, a scoring rule $S$ is proper when the following holds for any probabilistic forecast $F$ and true distribution $G$:
@@
Here $S(F,y)$ is the score of the predictive CDF at the observation $y$ and $\mathbb{E}_{y\sim G}S(F,y)$ denotes its expectation over the distribution of $G$. A scoring rule is thus called proper when in expectation it is minimal at the true distribution. It is called strictly proper if equality in equation 8 holds only when $F=G$.
@@
where $g$ is the density corresponding to $G$. Thus, when evaluating probabilistic forecasting models using scoring rules restricted to an outcome (such as events over a threshold) or scoring rules weighted by some weighting function (such as the HB-loss) this encourages hedging of the forecast distribution, meaning a forecaster is encouraged to deviate from the true distribution in their prediction by an amount determined by the weighting function. This also affects model training as can be seen in section 3.
@@
for a nonnegative weighting function $w$. This reduces to the unweighted CRPS (equation 4) when $w(x)\equiv 1$. The canonical weighting is indicator weighting: $w_{\tau}(x)=\mathbb{1}\{x\geq\tau\}$, when there is interest in values above (or below) a certain threshold $\tau$, which then leads to the so-called threshold-weighted CRPS (twCRPS):
@@
The weighted CRPS and especially the threshold-weighted versions have become frequently used tools for forecast evaluation of extreme events (e.g. Lerch and Thorarinsdottir, 2013; Baran and Lerch, 2015; Baran and Nemoda, 2016; Allen et al., 2021b). They have also been generalized to multivariate versions (Allen et al., 2023a, b). In the following we will focus on the twCRPS as given in equation 11.
@@
This is alluded to already in Allen et al. (2023a) and implemented in the scoringRules package (Allen, 2024). The proof is given in the Appendix. Thus intuitively when using the twCRPS a forecaster effectively uses the full CRPS to evaluate the distribution $\tilde{F}_{\tau}$ that has all probability mass below the threshold collapsed onto the threshold itself, whilst remaining the same above the threshold. That way the exact form of the distribution below $\tau$ only influences the score via the probability mass for exceeding or falling below the threshold.
@@
where $v_{\tau}=\max(y,\tau)$ as above. The formulation of the twCRPS as an expectation can be used to approximate it via samples from a predictive distribution, i.e. using the estimator
@@
Although the twCRPS has been frequently used for evaluation of probabilistic forecasting models for extremes, it has to this date not been used for training (post-processing) models in the literature or operations. However, given that weighted and threshold-weighted scores are proper – something that section 3 showed is important for model training – and the relevance of reliable forecasts for extremes, we will study their usage for the training of EMOS post-processing models for wind speed, for both truncated normal and truncated logistic EMOS models.
When comparing different models we will frequently use skill scores in the next section. Given forecasts $F_{1},\dots,F_{N}$, reference forecasts $F^{\text{ref}}_{1},\dots F^{\text{ref}}_{N}$ and observations $y_{1},\dots,y_{N}$ then a skill score in percent is given as:
$\text{Skill }(\%)=100\cdot\left(1-\frac{\sum_{i=1}^{N}\text{wCRPS}(F_{i},y_{i})}{\sum_{i=1}^{N}\text{wCRPS}(F^{\text{ref}}_{i},y_{i})}\right),$
for any CRPS weighting function $w$. It thus corresponds to the percentage improvement in the score over the reference forecasts.
@@ 5 Results: training with weighted scoring rules
We study the usage of the weighted CRPS for the training of post-processing models, to improve probabilistic predictions of extreme wind speeds. The weighted CRPS is a proper scoring rule, thus using it for evaluation and training will favour calibrated-sharp forecasts. We focus on the canonical threshold weighting due to the availability of closed-form expressions, which helps with the optimization. We consider training EMOS models using the following proper scores:
@@
Here the CRPS and log-likelihood models serve as baselines, whilst the twCRPS puts special emphasis on values above a threshold $\tau$. For optimizing the score functions we use the closed-form expressions for the CRPS and twCRPS for the truncated normal and truncated logistic distributions (see Appendix B). We also tried optimization using the sample twCRPS estimator (equation 15) and samples from the predictive distributions, but this proved slower and much less robust than optimization using the closed-form expressions.
@@ 5.1 Effect of weighted training
We consider thresholds $\tau$ defined as the location-wise 80th and 90th percentile of the training dataset and train using the scores in section 5. In practical applications most likely absolute thresholds defined by user needs will be of relevance, however, to show the general applicability of the method we use percentile based thresholds here as in Sharpe et al. (2018); Sharpe and Stewart (2019); Sharpe (2022). Table 1 in the Appendix shows the 80th and 90th percentile threshold in each cluster of stations. We report skill scores relative to the model trained using CRPS (“CRPS-trained model").
@@
Figure 3: Threshold-weighted CRPS skill score (left, a) and CRPS skill score (right, b), both relative to the CRPS-trained model, for wind speed EMOS using truncated normal and truncated logistic distributions. Models are trained using log likelihood, and twCRPS. Thresholds for training and evaluation of the twCRPS are defined as locationwise 80th and 90th percentile of the training dataset. The diamonds correspond to the results within each cluster and the bands to the spread.
@@
Figure 3a shows for all models the threshold-weighted CRPS skill score over the CRPS-trained model evaluated on the test set, for the 80th and 90th locationwise percentile thresholds (used for both training and evaluation) for both truncated normal and truncated logistic predictive distributions. Training using a twCRPS leads to improved predictions of extremes, as measured by the twCRPS on the test set. These improvements average around $0.8$% for the 80th and around $1.3$% for the 90th percentile threshold and up to $\sim 1.4$% and $\sim 2.5$% respectively. The twCRPS improvements also show in terms of Brier skill score (see Table 2 in the Appendix for the Brier skill score for both distributions for a variety of evaluation thresholds). The strongest improvements in terms of twCRPS are in cluster 1 corresponding to inland locations, followed by clusters 3 and 2 corresponding to coastal and more exposed inland locations, with cluster 4 (more exposed coastal locations) last. The log-likelihood training performs worse in terms of CRPS and twCRPS than the CRPS-trained model which is in line with previous studies (Gneiting et al., 2005; Gebetsberger et al., 2018).
@@
The improvements in terms of twCRPS however seem to be linked to a loss in CRPS as can be seen in Figure 3b which shows the CRPS skill over the CRPS-trained model, evaluated on the test set. The twCRPS-trained model loses CRPS skill, compared to the CRPS-trained model for both the 80th and even stronger for the 90th percentile threshold. Performance at the latter threshold improves more in terms of twCRPS, but also decreases more in terms of CRPS, irrespective of the predictive distribution. There seems to be a trade-off between CRPS and twCRPS performance. We investigate this more closely in Figure 4, which shows the twCRPS skill improvement against the CRPS skill loss for all clusters and both truncated normal and truncated logistic, for 80th and 90th percentile thresholds. There seems to be a near-linear association between CRPS skill loss and twCRPS skill improvement. Larger twCRPS improvements also seem to lead to larger CRPS losses, which are both associated with the larger of the two thresholds. In section 6 we will study ways of managing the trade-off between CRPS and twCRPS performance.
@@
Figure 4: CRPS skill loss against twCRPS skill gain, both relative to the CRPS-trained model, for all clusters, truncated normal (points) and truncated logistic (squares) predictive distribution and 80th (blue) and 90th (orange) percentile training and evaluation thresholds. The line corresponds to the line of best fit.
We also investigate the effect of twCRPS based training onto extremal calibration, focusing on the truncated normal distribution and 90th percentile threshold, although results for the other models look similar. Figure 5a shows QQ-plots of the conditional PIT values against quantiles from a uniform distribution for CRPS-, log-likelihood and twCRPS-based training. Under probabilistic calibration of the threshold exceedance distribution, these quantiles should be close to the diagonal. twCRPS-based training seems to improve extremal calibration, including in cluster 4 which has the most extreme wind speeds but in which only a small twCRPS score improvement can be found. A similar picture appears when considering the tail miscalibration TMCB (equation 5). This is shown in Figure 5b as a function of the evaluation threshold. For evaluation thresholds close to the 90th percentile training threshold the twCRPS-trained model performs similarly as the CRPS and log-likelihood trained one. However for higher thresholds it shows much better tail calibration. This is similar also for the 80th percentile model, which shows substantially improved calibration over the 90th percentile (not shown). The improvements in calibration thus seem to hold for a number of thresholds.
@@
In summary, by optimizing the twCRPS one can improve extreme event performance of forecasts, as measured by twCRPS, Brier Score and extremal calibration. This however comes at some cost in terms of CRPS – measuring the forecast performance for the whole distribution and the distribution body. The trade-off between CRPS and twCRPS performance can be seen as a classic distribution body-tail trade-off where an improved modelling of a distribution’s tails might come at the cost of capturing the distribution’s body and vice-versa. This seems to hold for training thresholds defined as 80th and 90th percentile of the distribution, however, in the next section we will study the effect of varying the training threshold.
Figure 5: Extremal calibration: a) By cluster: QQ plots of quantiles from a uniform distribution against of conditional PIT values for CRPS-, log-likelihood- and twCRPS-trained models, 90th percentile, truncated normal predictive distribution. b) TMCB as a function of the evaluation threshold for the CRPS-, log-likelihood and twCRPS-trained models, 90th percentile model, truncated normal predictive distribution. The diamonds represent the results in each cluster, whilst the line corresponds to the mean.
@@ 5.2 Effect of the training threshold
We study the effect of varying the threshold for training using the twCRPS and analyse this for the truncated normal case. Results for the truncated logistic distribution are similar. We train post-processing models using the twCRPS with thresholds as 60th, 80th, 90th and 95th percentile and evaluate against the twCRPS with a range of different thresholds between the 0th (corresponding to the unweighted CRPS) to 99th percentile of training dataset. Figure 6 shows by cluster the twCRPS skill (relative to the CRPS-trained model) as a function of the evaluation threshold (in m/s) for all four trained models. Results in terms of quantile can be found in Figure 14, and for the 90th percentile model a comparison against the raw ensemble can be seen in Figure 13, both in the Appendix. The model trained for a certain threshold generally seems to perform well at this threshold and at thresholds slightly lower and higher than the one it was trained for. All models experience a drop in performance near the 99th percentile, which generally corresponds to very high wind speed values which naturally will be very difficult to predict. Training with a twCRPS for a high quantile threshold also leads to skill loss at lower quantiles and it seems the higher the training threshold / quantile the worse the twCRPS skill loss for lower evaluation thresholds / quantiles also becomes, manifesting in increasing slope of the skill curves. This means that model performance can be improved for many thresholds by training with a twCRPS, but should the distribution body performance be important it might be sensible to use lower thresholds. In these cases however the skill improvements for high quantiles are also more modest. Some improvement can also be seen in terms of Brier score (see Figure 15 in the Appendix), at least for cluster 1, 2, and 3, although results are more varied.
@@
Figure 6: twCRPS skill score (over the CRPS-trained model) by cluster as a function of the evaluation threshold for twCRPS-trained models trained using different thresholds. Thresholds on the x-axis correspond to wind speeds between the 0th and 99th percentile of the training distribution. The vertical lines indicate the thresholds for which models were trained.
5.3 Effect of training datasize
We study the effect of the training datasize for training EMOS models using the twCRPS. For this we vary the number of stations that twCRPS-based EMOS models are trained on, allowing us to probe the amount of data needed. We define subsets of N stations by sampling at random with replacement from the dataset of 124 stations, ignoring the previous cluster assignments. EMOS models are then trained on the dataset of stations within this subset/cluster using CRPS and twCRPS, in order to calculate the twCRPS skill of the twCRPS-model on the subset-specific test set. This process is bootstrapped 50 times to get an indication of the uncertainty. We focus on the truncated normal predictive distribution for 80th and 90th percentile models and for models trained on the whole training data (one year and 9 months), as well as models trained on only one year of training data. Figure 7 visualizes the skill score as a function of the number of locations included in the training set.
The training datasize seems to have substantial influence on whether twCRPS-based training is successful. Consistent positive twCRPS skill on the test set can only be found when enough locations are included in each subset/cluster, for both percentiles and both the full and half training data. More locations are needed for the 90th percentile model than for the 80th percentile model and for the models using half the training data compared to the models trained on the full training data. Little benefit in terms of twCRPS skill can be seen for training on individual stations in this dataset, however this might be different if more training data per station would be available. Interestingly, even if there is no improvement in terms of twCRPS skill for smaller clusters (including locationwise training), there is generally still an improvement in terms of extremal calibration (not shown).
This dependence on training datasize might be due to multiple reasons. Extensive training data might be needed to infer tail properties well. Also, Lerch et al. (2017) report low discriminative power of the twCRPS in Diebold-Mariano test. Thus, in order for having adequate power to distinguish well between candidate EMOS models, large amounts of data might be needed. The training data requirements also seem to vary with the thresholds, with successful training of twCRPS models for larger thresholds needing more training data than for lower thresholds.
Figure 7: twCRPS skill score (over the CRPS-trained model) as a function of the number of locations included in the training set, for models trained on the full training dataset, as well as on only one year of training data. The bands represent 95% confidence intervals obtained through bootstrapping, whilst the lines correspond to the mean score. Truncated normal predictive distribution and 80th as well as 90th percentile thresholds.